# A template for the arxiv style

### v1

@@ 4.1 Control of false positives
In this subsection we define a null hypothesis for each subset $S\subseteq[D]$, which will help us decide if $S$ is a set of plausible causal parents:
$\tilde{H}_{0,S}:\begin{cases}&\text{there exists a distribution $F$ and $\gamma^{e}\in\mathbb{R}^{|S|}$, s.t. $\forall e\in[E]$:}\\ &\text{$Y^{e}=X^{e}_{S}\gamma^{e}+r^{e}$ and for all $i,j\in[n^{e}],i\neq j$:}\\ &\text{$r^{e}_{i}\sim F,r^{e}_{i}\perp\!\!\!\perp r^{e}_{j},r^{e}_{i}\perp\!\!\!\perp X^{e}_{S,i}$}\ .\end{cases}$
We note that $\tilde{H}_{0,S}$ corresponds essentially to Assumption 1 in relation to set $S$. With that we know in particular that, under Assumption 1, $\tilde{H}_{0,S^{*}}$ is true. While $\tilde{H}_{0,S}$ has an exact correspondence to Assumption 1, it is, however, not clear how to build a practical test based on it, that is also powerful against meaningful alternatives. For that reason we move towards a weaker but more practical formulation. Let
@@
where $A^{\dagger}$ denotes the generalized Moore-Penrose inverse of a matrix $A$ (Penrose, 1955), which always exists and is unique, and where we set $\tilde{\beta}^{e}_{S}=0$ for $S=\emptyset$. With this we can formulate a simpler null hypothesis as
$H_{0,S}:\begin{cases}&\text{there exists a distribution $F$ such that for all $e\in[E]$:}\\ &\text{$Y^{e}=X^{e}_{S}\tilde{\beta}^{e}_{S}+r^{e}$ and for all $i\in[n^{e}]:r^{e}_{i}\sim F$\ .}\end{cases}$
The following lemma, proven in Appendix A, establishes the relation between $H_{0,S}$ and $\tilde{H}_{0,S}$:
@@
Remark 1.
The definition of the null hypothesis ${H}_{0,S}$ already foreshadows the relevance of the different environments. Consider the case that we only have one environment $e_{1}$ and the rows of $X^{e_{1}}$ are i.i.d (independently and identically distributed) samples from the same distribution. In that case even $S=\emptyset$ fulfills ${H}_{0,S}$ for $\gamma^{e_{1}}$ being the null-vector and $r^{e}=\beta^{e_{1}}X^{e_{1}}+\varepsilon^{e_{1}}$. This is in accordance with our intuition of thinking about the environments as interventions. If we only have a single intervention, but no control, we cannot draw causal conclusions about the effect of that intervention.
Based on the hypothesis $H_{0,S}$ we are going to consider the following meta-procedure: Recall that $(\boldsymbol{X},\boldsymbol{Y})=\{(X^{e},Y^{e})\}_{e\in[E]}$ are our observations, and let $\phi_{S}(\boldsymbol{X},\boldsymbol{Y})\in\{0,1\}$ be a test function, such that $\phi_{S}(\boldsymbol{X},\boldsymbol{Y})=1$ indicates we reject ${H}_{0,S}$. We know that $\tilde{H}_{0,S^{*}}$ holds by Assumption 1, and thus Lemma 1 implies $H_{0,S^{*}}$ also holds, and we expect that with high probability $\phi_{S^{*}}(\boldsymbol{X},\boldsymbol{Y})=0$. With this in mind we view all $S$ for which $\phi_{S}(\boldsymbol{X},\boldsymbol{Y})=0$ as a plausible set, and naturally define the estimator $\tilde{S}$ of $S^{*}$ as
$\tilde{S}\vcentcolon=\bigcap_{S:\phi_{S}(\boldsymbol{X},\boldsymbol{Y})=0}S.$
Based on the above considerations we have the following result:
@@
Let $\alpha\in(0,1)$. Consider a class of test functions $\phi_{S}$ for all $S\subseteq[D]$. Assume this class of tests satisfies $P[\phi_{S^{*}}(\boldsymbol{X},\boldsymbol{Y})=1\mid{H}_{0,S^{*}}\text{\ \ holds}]\leq\alpha$. Then under Assumption 1 we have $\tilde{S}\subseteq S^{*}$ with probability of at least $1-\alpha$.
This result, which is essentially Theorem 1 from Peters et al. (2016), is a simple consequence of the fact that $P[\phi_{S^{*}}(\boldsymbol{X},\boldsymbol{Y})=1\mid{H}_{0,S^{*}}\text{\ \ holds}]\leq\alpha$. Note that for this result it is only important to guarantees the good behavior of the testing procedure for the true set $S^{*}$.
@@ 4.2 Control of false negatives
The previous section shows that we can guarantee that $\tilde{S}\subseteq S^{*}$ with arbitrarily high probability. However, such a guarantee can be trivially obtained for the choice $\tilde{S}=\emptyset$. A natural question to ask is when is it possible to have a class of tests that ensure that $\tilde{S}=S^{*}$ with high probability. Naturally the answer to this question is more intricate, and depends crucially on how much information is present on data. For the following result we focus on scenarios where one has an arbitrarily large amount of data in each environment, and identify conditions under which the methodology above might be able to identify $S^{*}$ as the only possible set of plausible causal parents.
Specifically, suppose one has access to $\tilde{\beta}^{e}_{S}$ for any $e\in[E]$ and $S\subseteq[D]$. In practical terms this would essentially require to observe the randomly generated data in all environments an infinite number of times. Given access to $\tilde{\beta}^{e}_{S}$, which requirements on the heterogeneity of the environments are needed to ensure one can pinpoint $S^{*}$ as the only set of plausible causal parents? Within a fairly general class of structural equation models for the covariates $X^{e}$, heterogeneity is almost always guaranteed through the changing parameters $\beta^{e}$, and only in rather pathological combinations of parameters problems will arise. This is made precise in Theorem 1. The class of structural equation models for which we can show identifiability is given through the following assumption:
@@
$X^{e}_{d}\vcentcolon=f_{d}(X^{e}_{\textbf{PA}(d)})+\delta^{e}_{d},$
where $f_{d}$ is a polynomial of finite degree in $X^{e}_{y}=Y^{e}$ if $y\in\textbf{PA}(d)$, but otherwise arbitrary. Here $\delta^{e}_{d}=(\delta^{e}_{d,1},\cdots,\delta^{e}_{d,n^{e}})$ is a random noise vector such that
$\delta^{e}_{d}\sim\mathcal{D}^{e}_{d}$
$\forall:i\in[n^{e}]\quad\delta^{e}_{d,i}\perp\!\!\!\perp X^{e}_{u,i}\quad\text{\ for all \ }u\in\textbf{NDE}(d),$ (4)
@@
Informally speaking the noise terms $\delta_{d}^{e}$ ensure that each covariate introduces unique information and prevent that the causal parents $X^{e}_{S^{*}}$ lie in the column space of any other subset of variables $X^{e}_{S}$ with $S^{*}\not\subseteq S$. The need for the explicit variances $\Delta^{e}_{d,i}$ is due to our proof technique, and we contextualize this assumption further after we state the theorem. The polynomial dependence on $Y^{e}$ simplifies our proof, could, however, be replaced by other regularity assumptions on $f_{d}$. With this in hand we are ready to state our main identifiability result:
@@
Let $S\subset[D]$ such that $S^{*}\not\subseteq S$ and take any two environments $e,v\in[E]$ with $e\neq v$, such that environment $e$ fulfills the data generation mechanism from Assumption 2 with variances $\Delta^{e}$. Suppose Assumption 1 holds with parameters $\beta^{e},\beta^{v}$. Then there is a set $M_{0}\subset\mathbb{R}^{|S^{*}|\times|S^{*}|}\times(0,\infty)^{D\times n^{e}}$ with Lebesgue measure zero, such that if
$(\beta^{v},\beta^{e},\Delta^{e})\in\mathbb{R}^{|S^{*}|\times|S^{*}|}\times(0,\infty)^{D\times n^{e}}\setminus M_{0}$

### v2

@@ 4.1 Control of false positives
Given a set $S\subseteq[D]$ consider the following null hypothesis:
$\tilde{H}_{0,S}:\begin{cases}&\text{$\exists$ a distribution $F$ and $\gamma^{e}\in\mathbb{R}^{|S|}$, s.t.$\forall e\in[E]:$ }\\ &\text{$Y^{e}=X^{e}_{S}\gamma^{e}+r^{e}$ and $\forall i,j\in[n^{e}],i\neq j$:}\\ &\text{$r^{e}_{i}\sim F,r^{e}_{i}\perp\!\!\!\perp r^{e}_{j},r^{e}_{i}\perp\!\!\!\perp X^{e}_{S,i}$}\ .\end{cases}$
We note that $\tilde{H}_{0,S}$ corresponds to Assumption 1 when $S^{*}$ is replaced by $S$, which in particular implies that under Assumption 1 $\tilde{H}_{0,S^{*}}$ is true. However, it is not clear how to build a practical test based on $\tilde{H}_{0,S}$ that is also powerful against meaningful alternatives. For that reason we move towards a weaker but more practical formulation. Let
@@
where $A^{\dagger}$ denotes the generalized Moore-Penrose inverse of a matrix $A$ (Penrose, 1955), with the convention that $\tilde{\beta}^{e}_{\emptyset}=0$. Formulate the following relaxation of $\tilde{H}_{0,S}$:
$H_{0,S}:\begin{cases}&\text{$\exists$ a distribution $F$ such that for all $e\in[E]$:}\\ &\text{$Y^{e}=X^{e}_{S}\tilde{\beta}^{e}_{S}+r^{e}$ and $\forall i\in[n^{e}]:r^{e}_{i}\sim F$\ .}\end{cases}$
The following lemma, which is proven in Appendix A.1, establishes the relation between $H_{0,S}$ and $\tilde{H}_{0,S}$:
@@
Suppose we have access to a collection of tests corresponding to the above null hypothesis $H_{0,S}$. Specifically, given the observations $(\boldsymbol{X},\boldsymbol{Y})$ let $\phi_{S}(\boldsymbol{X},\boldsymbol{Y})\in\{0,1\}$ be a test function, such that $\phi_{S}(\boldsymbol{X},\boldsymbol{Y})=1$ indicates we reject ${H}_{0,S}$. We know that $\tilde{H}_{0,S^{*}}$ holds by Assumption 1, and thus Lemma 1 implies $H_{0,S^{*}}$ also holds, and we expect that with high probability $\phi_{S^{*}}(\boldsymbol{X},\boldsymbol{Y})=0$. With this in mind we view all $S$ for which $\phi_{S}(\boldsymbol{X},\boldsymbol{Y})=0$ as a plausible set, and naturally define the estimator $\tilde{S}$ of $S^{*}$ as
$\tilde{S}\vcentcolon=\bigcap_{S:\phi_{S}(\boldsymbol{X},\boldsymbol{Y})=0}S.$ (4)
This definition of the parent estimator $\tilde{S}$ ensures control over false discoveries:
@@
Let $\alpha\in(0,1)$. Consider a class of test functions $\phi_{S}$ for all $S\subseteq[D]$ that satisfies $P[\phi_{S^{*}}(\boldsymbol{X},\boldsymbol{Y})=1\mid{H}_{0,S^{*}}\text{\ \ holds}]\leq\alpha$. Then we have $\tilde{S}\subseteq S^{*}$ with probability of at least $1-\alpha$.
This result, which is essentially Theorem 1 from Peters et al. (2016), is a simple consequence of the fact that $P[\phi_{S^{*}}(\boldsymbol{X},\boldsymbol{Y})=1\mid{H}_{0,S^{*}}\text{\ \ holds}]\leq\alpha$. Note that for this result it suffices to guarantee good behavior of the testing procedure for the true set $S^{*}$.
@@ 4.2 Control of false negatives
Proposition 1 guarantees $\tilde{S}\subseteq S^{*}$ with arbitrarily high probability. In other words, we are guaranteed to not include non-causal parents with high probability. However, that can be trivially obtained for the choice $\tilde{S}=\emptyset$. Naturally, we ask under which assumptions one can have a class of tests that ensure that $\tilde{S}=S^{*}$ with high probability as well. The answer to this question is significantly more intricate, and depends crucially on how much information is present in the data. To shed some light on this matter we consider a population setting, effectively focusing on scenarios where one has an arbitrarily large amount of data in each environment.
Specifically, suppose one has access to $\tilde{\beta}^{e}_{S}$ for any $e\in[E]$ and $S\subseteq[D]$. Theorem 1 below shows that, within a fairly general class of structural equation models for the covariates $X^{e}$, the proposed approach can identify $S^{*}$, and only in rather pathological combinations of parameters issues might arise. The class of structural equation models for which we can show identifiability is given through the following assumption:
@@
$X^{e}_{d}\vcentcolon=f^{e}_{d}(X^{e}_{\textbf{PA}(d)})+\delta^{e}_{d},$
where $f^{e}_{d}$ is a polynomial of finite degree in $X^{e}_{y}=Y^{e}$ if $y\in\textbf{PA}(d)$, but otherwise arbitrary. More specifically in that case $f^{e}_{d}$ has the form
$f^{e}_{d}(X^{e}_{\textbf{PA}(d)})=\sum_{k=0}^{K}(Y^{e})^{k}g^{e}_{k}\left(X^{e}_{\textbf{PA}(d)\setminus y}\right),$
where $K<\infty$ and the functions $g_{k}^{e}$ are arbitrary. In the above $\delta^{e}_{d}=(\delta^{e}_{d,1},\cdots,\delta^{e}_{d,n^{e}})\sim\mathcal{D}^{e}_{d}$ is a random noise vector such that
$\forall:i\in[n^{e}],u\in\textbf{NDE}(d):\quad\delta^{e}_{d,i}\perp\!\!\!\perp X^{e}_{u,i}$ (5)
@@
Informally speaking the noise terms $\delta_{d}^{e}$ ensure that each covariate introduces unique information and prevent that the causal parents $X^{e}_{S^{*}}$ lie in the column space of any other subset of variables $X^{e}_{S}$ with $S^{*}\not\subseteq S$. The explicit variances $\Delta^{e}_{d,i}$ are needed due to our proof technique, but we highlight that we do not require $\Delta^{e}_{d,i}$ to be heterogeneous in the environments. The polynomial dependence on $Y^{e}$ simplifies our proof, could, however, be replaced by other regularity assumptions on $f^{e}_{d}$. With this in hand we are ready to state our main identifiability result:
@@
Let $S\subset[D]$ such that $S^{*}\not\subseteq S$ and take any two environments $e,v\in[E]$ with $e\neq v$, such that environment $e$ fulfills the data generation mechanism from Assumption 2 with variances $\Delta^{e}$. Suppose Assumption 1 holds with parameters $\beta^{e},\beta^{v}$. Then there exists a set $M_{0}\subset\mathbb{R}^{|S^{*}|\times|S^{*}|}\times(0,\infty)^{D\times n^{e}}$ with Lebesgue measure zero, such that if
$(\beta^{v},\beta^{e},\Delta^{e})\not\in M_{0}$